OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below. If a = x and x = b, then a = b

OpenStudy (anonymous):

Transitive Property

OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below. 3(x + y) = 3x + 3y

OpenStudy (anonymous):

distributive property

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below. 2 + x + y = x + 2 + y

OpenStudy (anonymous):

nvm